OpenStudy (anonymous):

y=0.5cos2/3(x+pi)-2 How to graph this? i'm really stuck :'( please help!

OpenStudy (unklerhaukus):

\[y=\frac12\cos\left(\frac23(x+\pi)\right)-2 \]

OpenStudy (unklerhaukus):

can you draw cos(x)? the 1/2 out the front will mean the amplitude is half the 2/3 means wave is squashed a bit in the x direction the pi puts the graph out of phase by half a wavelength the -2 at the end means the whole thing is shifted down 2 units

OpenStudy (unklerhaukus):

\[4+2y=\sin\left(\frac{2π}{3}\right)\]

OpenStudy (anonymous):

okay thanks ^_^ i'm not sure what the units for the x-axis will be?

OpenStudy (unklerhaukus):

**\[4+2y=\sin\left(\frac{2x}{3}\right)\]

OpenStudy (unklerhaukus):

i would use pi radiens as the units for the x axis

OpenStudy (anonymous):

okay thanks so much! i really do appreciate it!! :D